Sin2x+sin2y=1/2 , cos2x+cos2y=3/2 then the value of cos.cos (x-y) is?

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Using the identity sinA + sinB
we get sin(x+y)cos(x-y) = ¼ and cos(x+y)cos(x-y) = ¾.
Squaring and adding both we get cos(x-y) = sqrt(10)/4.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)